Theo needs some ribbon to go around 5 rectangular cakes. Each cake is 28 cm long and 32 cm wide. How much ribbon will he need?  

The ribbon costs $1 per metre. How much will the ribbon cost Theo? 


1
Expert's answer
2021-10-19T20:13:03-0400

For one cake Theo needs: "(28+32)*2=120" cm.

For five cakes Theo needs: "120*5=600" cm."=6" m.

The cost of ribbon: "1*6=\$6".